Reports 1-1 of 1 Clear search Modify search
DGS (General)
shoichi.oshino - 12:39 Friday 29 September 2023 (26975) Print this report
Cleaning network optical fiber
As requested by Ikeda-san, I did network optical fiber cleaning.
This is one of the countermeasures for the IPC glitch problem.
 
[Activity log]
1057 k1dc0 PC side in
1058 k1dc0 PC side out
1059 k1dc0 SW side in
1100 k1dc0 SW side out
1101 k1nds0 PC side
1102 k1nds0 SW side
1104 k1nds1 PC side
1105 k1nds1 SW side
1106 k1fw0 PC side in
1107 k1fw0 SW side in
1108 k1fw1 PC side in
1109 k1fw1 SW side in
1111 k1tw0 PC side
1111 k1tw0 SW side
1113 k1tw1 PC side
1113 k1tw1 SW side
1114 k1bcst0 SW side
1116 k1bcst0 PC side
1118 finished
1121 Requested DAQ_KILL
Comments to this report:
shoichi.oshino - 12:40 Friday 29 September 2023 (26976) Print this report
Yamamoto-san noticed me that he could not see the data.
First, I checked if the frame file was generated.
The latest frame files have been created for both DAQ systems.
Next, I checked to see if the directory was mounted on NDS servers.
I can access each frame file directory from NDS servers.
After that, I requested DAQ_KILL from the guardian again.
After the DAQ was restored, I confirmed that I could view the data with ndscope. 
satoru.ikeda - 18:19 Friday 06 October 2023 (27113) Print this report

We checked for a week and the time of occurrence of the IPC glitch changed to around 20:00, but there was no effect on the presence or absence of the occurrence.

Non-image files attached to this comment
Search Help
×

Warning

×